# Dual-Level Boost Network for Long-Tail Prohibited Items Detection in X-ray Security Inspection

### v1

@@ Abstract
The detection of prohibited items in X-ray security inspections is vital for ensuring public safety. However, the long-tail distribution of item categories, where certain prohibited items are far less common, poses a big challenge for detection models, as rare categories often lack sufficient training data. Existing methods struggle to classify these rare items accurately due to this imbalance. In this paper, we propose a Dual-level Boost Network (DBNet) specifically designed to overcome these challenges in X-ray security screening. Our approach introduces two key innovations: (1) a specific data augmentation strategy employing Poisson blending, inspired by the characteristics of X-ray images, to generate realistic synthetic instances of rare items which can effectively mitigate data imbalance; and (2) a context-aware feature enhancement module that captures the spatial and semantic interactions between objects and their surroundings, enhancing classification accuracy for underrepresented categories. Extensive experimental results demonstrate that DBNet improves detection performance for tail categories, outperforming sota methods in X-ray security inspection scenarios by a large margin (+17.2%), thereby ensuring enhanced public safety [^1] Code will be released upon acceptance..
@@ 1 Introduction
As crowd density increases in public transportation hubs [2, 38], ensuring public safety through effective security inspections has become more critical than ever. X-ray scanners [43], widely deployed to inspect luggage and produce complex imagery, play a crucial role in these inspections. However, the accuracy of identifying prohibited items can be compromised due to the limitations of manual inspection, particularly under conditions of prolonged concentration, which may lead to missed detention of prohibited items and potential security breaches. Therefore, there is a pressing need for an accurate and automated detection to support these safety efforts. Advances in deep learning [16, 8, 15], particularly CNNs, offer promising solutions by treating AI inspection as an object detection task [49, 48, 24] in computer vision community [32].
Figure 1: Performance drop of head and tail classes. DBNet addresses this issue with both data-level and model-level strategies.
Recent researchers have focused extensively on improving detection accuracy for a diverse set of prohibited items. On the data front, a number of X-ray image benchmarks have been released, including GDXray [21], SIXray [23], and OPIXray [42]. On the model front, traditional CNN-based approaches [44, 10, 25] trained on generic object detection datasets have often struggled to achieve satisfactory results in this unique context. Unlike natural images, which often contain simple and distinctive visual information [6, 29], X-ray images are characterized by a lack of strong discriminative properties and often contain heavy visual noise [37, 22, 20]. Significant strides have been made in advancing the detection of prohibited items in X-ray imagery, with a lot of breakthroughs in model performance.
However, a critical challenge remains largely unaddressed by most existing approaches: the long-tail distribution of item appearances in real-world scenarios. Common items, such as knives, are disproportionately represented in training datasets, whereas rare items, like improvised explosive devices or unusual contraband, are far less frequent. This data imbalance causes the performance drop of detection models and prevents these models fro m identifying infrequent prohibited items accurately, which poses a significant risk in real-world practical applications.
X-ray images have unique characteristics that make detecting items more challenging. Unlike natural images, where features like color and texture help with identification, X-ray images primarily show the internal structure of objects, removing many obvious visual clues. Furthermore, items inside luggage often overlap and block each other, making it difficult to clearly identify individual objects. This means that detecting rare items in X-ray images is quite different from doing so in regular, visible-light images. In the absence of distinctive attributes such as color and texture, combined with frequent occlusions, the accurate recognition and classification of uncommon items become increasingly more challenging. As Figure 2 illustrated, the primary challenge lies not only in location but also in accurately classifying the items of tail categories.
To address these challenges, we propose DBNet (Figure 1), a dual-level boost framework that is specifically designed to enhance the detection of tail prohibited items at both the data and model levels. To be specific, DBNet integrates a novel data augmentation approach using Poisson blending, which generates realistic synthetic instances of rare items by cropping them from their original images and seamlessly integrating them into other images. On the model level, DBNet incorporates a context-aware feature enhancement module that explicitly captures interactions between objects and their surroundings. By utilizing spatial and semantic relationships between target items and their context, this module allows the model to interpret complex scenes more effectively. Together, these two augmentation techniques form a comprehensive solution that significantly improves detection accuracy for low-frequency classes, offering a robust framework to address long-tail distribution challenges in detecting prohibited items in X-ray imagery.
@@
Figure 2: Preliminary experimental visualization. The left side shows the ground truth, and the right displays the predictions. We observe that objects from tail classes are often misclassified as other categories. For example, when a hammer is present, its parts should not be misclassified as a knife. This highlights the motivation for incorporating surrounding object context features.
Extensive experimental evaluation demonstrates that DBNet outperforms existing sota methods by a large margin, particularly for tail prohibited items. Our approach not only improves detection accuracy but also enhances model robustness in real-world X-ray security inspections. To sum up, this work aims to offer a comprehensive solution to the challenge of long-tail distribution in X-ray security screening, ultimately contributing to more reliable automated security systems. Our main contributions are as follows:
  • We propose a novel data augmentation technique using Poisson blending to generate realistic instances of rare hazardous materials, effectively mitigating class imbalance in X-ray imagery and boosting the performance.
  • We introduce a context-aware feature module that captures spatial and semantic relationships, crucial for detecting prohibited items that are often occluded in X-ray images, thus improving the detection performance.
@@
  • Extensive experiments demonstrate that DBNet outperforms existing methods by a large margin (+17.2%), offering a robust solution for accurate detection of tail prohibited items in real-world X-ray security inspection.
Figure 3: Framework overview of the proposed DBNet for long-tail X-ray prohibited items detection. The framework consists of two key components: data-level Poisson blending augmentation and model-level related proposals aggregation. These components cooperate to address the challenges of imbalanced data distribution, enhancing the detection accuracy for tail classes in complex X-ray imagery.
@@ 2.2 Long-tail Vision Tasks
The challenge of long-tail visual recognition has gained considerable attention due to the imbalance between the frequency of common (head) classes and rare (tail) classes in many real-world datasets. Various approaches have been proposed to address the inherent difficulty in learning effective representations for tail classes in the presence of data imbalance. Class rebalancing strategies, such as re-sampling and re-weighting, have been extensively studied to mitigate the bias towards head classes. For instance, [7] introduced the concept of an effective number of samples to create better class-balanced loss weights, resulting in improved performance for underrepresented classes. Another popular strategy [4] involves using data augmentation or synthetic data generation techniques to artificially increase the number of instances for tail classes, thereby balancing the class distribution. These re-sampling and augmentation techniques are effective in addressing class imbalance but can introduce issues like overfitting or under-representing the true diversity within each class.
Another line of research focuses on improving feature learning in long-tail recognition through meta-learning and contrastive learning approaches. [14] proposed Decoupled Training, which first learns a backbone feature extractor using a conventional cross-entropy loss and then fine-tunes the classifier with a rebalanced scheme to handle the data imbalance more effectively. [1] introduced Label-Distribution-Aware Margin (LDAM) Loss, which explicitly modifies the decision boundary for each class based on its frequency, enforcing larger margins for tail classes. In addition, contrastive learning has emerged as a powerful way to learn more generalized feature representations that are robust to class imbalance, such as the work by [33], who proposed a contrastive feature distribution alignment to ensure feature consistency between head and tail classes.
@@ 3 Method
In this section, we elaborate on the details of the proposed Dual-level Boost Network (DBNet). We begin with an analysis of the long-tail X-ray prohibited items detection task setting in Section 3.1. Next, we introduce the framework of DBNet in Section 3.2. We then describe the data-level augmentation and model-level augmentation in Sections 3.3 & 3.4, respectively. Finally, Section 3.5 provides a detailed illustration of the network training process.
3.1 Task Overview
In the proposed long-tail X-ray prohibited items detection task, the goal is to maximize detection performance on tail classes. The tail classes represent those categories with insufficient number of training instances. This goal can be formally expressed through a set of formulations to make this objective explicit. Formally, the dataset for this long-tail detection task can be defined as follows: Let $C=\{c_{1},c_{2},\dots,c_{m}\}$ represent the set of all object categories, where $m$ is the total number of classes. The classes are partitioned into head classes $C_{head}$ and tail classes $C_{tail}$:
$C_{head}\cup C_{tail}=C$ (1)
$C_{head}\cap C_{tail}=\varnothing$ (2)
The number of instances for head classes is much larger than the one for tail classes, i.e., for each head class $c_{i}\in C_{head}$, the number of training samples $N_{c_{i}}$ is far greater than the number of samples $N_{c_{j}}$ for a tail class $c_{j}\in C_{tail}$.
To formalize, we define a dataset consisting of a set of training samples $\{(I_{k},B_{k},C_{k})\}_{k=1}^{K}$, where $I_{k}$ represents an input image, $B_{k}=\{b_{1}^{k},b_{2}^{k},\dots,b_{n_{k}}^{k}\}$ denotes the set of boxes, and $C_{k}=\{c_{1}^{k},c_{2}^{k},\dots,c_{n_{k}}^{k}\}$ represents the class labels in image $I_{k}$. The long-tail distribution is defined as:
$N_{c_{head}}\gg N_{c_{tail}},\forall c_{head}\in C_{head},\forall c_{tail}\in C_{tail}$ (3)
where $N_{c_{head}}$ and $N_{c_{tail}}$ represent the number of instances for head and tail classes, respectively. This imbalance poses challenges in training, as standard training processes may be biased toward the head classes, resulting in poor generalization for tail classes.
The primary goal is to maximize the detection performance for tail classes, which can be expressed in terms of maximizing the detection metrics such as mean Average Precision (mAP) for the tail classes. We denote the mAP for the tail classes as $\text{mAP}_{tail}$, which is calculated over all tail classes $C_{tail}$. The objective can be formulated as $\max\,\text{mAP}_{tail}$, where:
$\text{mAP}_{tail}=\frac{1}{|C_{tail}|}\sum_{c\in C_{tail}}\text{AP}_{c}$ (4)
where $\text{AP}_{c}$ represents the Average Precision for class $c$. Our objective is to directly improve performance over this metric, focusing on better detection specifically for tail classes.
@@ 3.2 Framework Overview
Figure 3 provides an overview of the proposed framework. The architecture of DBNet consists of two distinct pipelines: data-level augmentation and model-level augmentation, both of which aim to improve detection performance in scenarios involving imbalanced distributions.
@@
In the data-level augmentation pipeline, a core strategy involves addressing the severe class imbalance, particularly between the head and tail classes. Instances belonging to tail classes are cropped from their original images and treated as foreground objects. These cropped instances are then seamlessly inserted into other images using Poisson image editing techniques, which ensure a natural and coherent integration by maintaining consistent lighting and blending with the background. This augmentation method enhances the presence of underrepresented classes, thereby increasing the diversity of training examples and providing a more balanced distribution of instances across different categories. The main purpose of this technique is to alleviate the imbalance in the dataset by effectively creating synthetic samples for the tail classes, which can significantly improve the model’s ability to generalize across tail classes.
The model-level related proposals aggregation pipeline takes a different approach by focusing on the contextual relationships between objects in an image. Specifically, this component aims to utilize the spatial and semantic correlations between the target object and the surrounding items. By incorporating contextual information during the detection process, the model can better interpret complex scenes and refine its classification decisions. For instance, the presence of certain objects in the vicinity of the target may increase the likelihood of a particular classification outcome, considering that prohibited items are often surrounded by related items, leveraging these spatial dependencies can improve classification accuracy. This pipeline emphasizes that objects are not detected in isolation; rather, their surrounding context provides crucial clues that can aid in distinguishing between similar categories and reducing misclassification. By capturing these contextual dependencies, the proposed framework aims to improve the robustness of the model, particularly in challenging detection scenarios with occlusions or cluttered backgrounds.
Together, these two strategies form a comprehensive framework for enhancing detection performance in long-tail object recognition tasks. The data-level augmentation addresses the issue of data scarcity and imbalance by augmenting tail class instances, whereas the related proposals aggregation focuses on leveraging contextual relationships to provide richer and more nuanced feature representations. By integrating both strategies, DBNet aims to significantly enhance the detection accuracy for underrepresented classes, ensuring a balanced and context-aware detection process that is capable of handling the intricacies.
3.3 Data-level Poisson Blending Augmentation
The data-level augmentation component in the proposed DBNet aims to address the imbalance between the head and tail classes by enhancing the representation of tail classes through suitable data augmentation. We formally describe the data augmentation used to mitigate this class imbalance.
@@ 3.3 Data-level Poisson Blending Augmentation
Consider a dataset $D=\{(I_{k},B_{k},C_{k})\}_{k=1}^{K}$, where each image $I_{k}$ contains a set of bounding boxes $B_{k}=\{b_{1}^{k},b_{2}^{k},\dots,b_{n_{k}}^{k}\}$ and corresponding class labels $C_{k}=\{c_{1}^{k},c_{2}^{k},\dots,c_{n_{k}}^{k}\}$. In a typical long-tail distribution, let the classes be denoted by $C=C_{head}\cup C_{tail}$, with $C_{head}$ representing the frequently occurring head classes and $C_{tail}$ representing the rarely occurring tail classes. The number of instances $N_{c}$ for a tail class $c\in C_{tail}$ is significantly smaller than for any head class $c^{\prime}\in C_{head}$, i.e., $N_{c}\ll N_{c^{\prime}}$. This severe class imbalance reduces the ability of the model to learn effective features for tail classes.
To alleviate this issue, we propose a data-level augmentation technique that involves instance cropping and Poisson fusion to generate synthetic data that increases the occurrence of tail class instances. The top-left part of Figure 3 illustrates the detailed implementation process. Formally, for each tail class $c\in C_{tail}$, we perform the following operations: Instance Cropping. For every image $I_{k}$ containing an instance of class $c$, the bounding box $b_{i}^{k}\in B_{k}$ that corresponds to class $c$ is cropped to create a sub-image $I_{c,i}=I_{k}[b_{i}^{k}]$. The set of cropped instances is denoted by $S_{c}=\{I_{c,i}\}$ for all $b_{i}^{k}$ corresponding to $c$. In the actual execution process, the existing bounding boxes do not precisely enclose the foreground objects and often include mixed background information. To address this issue, we manually annotated segmentation masks that contain only the foreground instances, allowing for accurate cropping of the foreground images. Poisson Fusion. Each cropped instance $I_{c,i}\in S_{c}$ is inserted into a different background image $I_{bg}$ using Poisson image editing to create a natural blend. New composite image $I_{new}$ can be represented as:
$I_{new}=\text{PoissonFusion}(I_{bg},I_{c,i},b_{new})$ (5)
where $b_{new}$ represents the new bounding box coordinates of the inserted object within $I_{bg}$. Poisson fusion ensures that the boundaries of the inserted object match the illumination and texture of the new background, resulting in a visually coherent image. Augmented Dataset. The newly generated images $I_{new}$, along with their corresponding updated bounding boxes and labels, are added back to the training dataset. Thus, the new augmented dataset becomes:
$D^{\prime}=D\cup\{(I_{new},B_{new},C_{new})\}$ (6)
where $B_{new}$ and $C_{new}$ are the augmented bounding boxes and labels, respectively.
@@
By applying these augmentation techniques to tail class instances, the effective number of training samples for augmented class $c\in C_{tail}$ is increased, thereby reducing the overall class imbalance. The augmentation operation aims to balance the effective sample size across classes, improving ability to generalize to tail classes during training.
3.4 Model-level Related Proposals Aggregation
In real-world scenarios, metallic tools have simple shapes, making them prone to confusion. Figure 2 illustrates a typical case of misclassification. The model-level related proposals aggregation strategy aims to enhance the model’s capacity to learn discriminative features for both head and tail classes by utilizing contextual information from surrounding objects. Unlike data-level augmentation, which focuses on balancing the number of training samples, model-level augmentation improves the feature representation by incorporating spatial relationships among objects. The detailed process is as follows:
To begin, for proposals generated by detection heads, each is associated with a confidence score indicating the likelihood that it contains an object. Let the initial set of proposals be $\{p_{1},p_{2},\dots,p_{N}\}$ with corresponding confidence scores $\{s_{1},s_{2},\dots,s_{N}\}$. We select the top $k$ proposals based on these confidence scores:
$P_{\text{top-k}}=\{p_{i_{1}},p_{i_{2}},\dots,p_{i_{k}}\}$ (7)
where $s_{i_{1}}\geq s_{i_{2}}\geq\dots\geq s_{i_{k}}$. Higher-confidence proposals represent more prominent prohibited items in the whole image, providing reliable features for relational reasoning.

### v2

# Enhancing Prohibited Item Detection through X-ray-Specific Augmentation and Contextual Feature Integration

@@ Abstract
X-ray prohibited item detection faces challenges due to the long-tail distribution and unique characteristics of X-ray imaging. Traditional data augmentation strategies, such as copy-paste and mixup, are ineffective at improving the detection of rare items due to the complex interactions between overlapping objects. Furthermore, X-ray imaging removes easily distinguishable features like color and texture, making it difficult to differentiate between visually similar categories. To address these challenges, in this work, we propose the X-ray Imaging-driven Detection Network (XIDNet). Inspired by the unique characteristics of X-ray imaging, this network introduces two key innovations: a novel X-ray-specific augmentation strategy that generates more realistic training samples for rare items, thereby improving detection performance for categories with insufficient samples, and an contextual feature integration algorithm that captures the spatial and semantic interactions between objects and surroundings under X-ray imaging, enhancing the model’s ability to distinguish between similar categories. Extensive experimental results show that XIDNet effectively leverages X-ray imaging characteristics to significantly improve detection performance, outperforming popular SoTA methods by up to +17.2% in tail categories.
@@ 1 Introduction
As crowd density increases in public transportation hubs [2, 35], ensuring public safety through effective security inspections has become more critical than ever. X-ray scanners [40], widely deployed to inspect luggage and produce complex imagery, play a crucial role in these inspections. However, the accuracy of identifying prohibited items can be compromised due to the limitations of manual inspection, particularly under conditions of prolonged concentration, which may lead to missed detention of prohibited items and potential security breaches. Therefore, there is a pressing need for an accurate and automated detection to support these safety efforts. Advances in deep learning [18, 16], particularly CNNs, offer promising solutions by treating AI inspection as an object detection task [49, 48] in computer vision community [30].
Figure 1: Comparison of existing data augmentation methods for common and X-ray images. Traditional methods like copy-paste and mixup fail to capture X-ray-specific characteristics, leading to a noticeable gap between synthetic and real images.
Recent research has made significant progress in improving the detection accuracy of prohibited items in X-ray images. Several X-ray benchmarks, such as GDXray [22], SIXray [23], and OPIXray [39], have been introduced to provide valuable resources for this task. In real-world scenarios, the frequency of different categories varies, leading to a long-tail distribution in prohibited items detection data, as seen in datasets like HiXray [32]. On the model front, traditional CNN-based approaches [41, 9] have demonstrated promising results in X-ray object detection, leveraging advancements in deep learning and object recognition. Despite the challenges unique to X-ray imagery, such as noise and the absence of strong visual cues like color and texture, these methods have pushed the boundaries of detection performance, showing significant improvements in identifying a wide range of prohibited items.
However, despite these advancements, existing methods still face substantial limitations: (1) The unique imaging mechanism of X-rays often results in complex coupling between overlapping objects, making it challenging for traditional data augmentation strategies, such as copy-paste [11] and mixup [42], to effectively capture relationships (see Fig. 1). Moreover, these techniques may distort the environmental context of the X-ray original image, negatively impacting performance, particularly for tail categories. (2) Color filtering by X-ray imaging erases key distinguishing features, making it difficult to differentiate between visually similar objects (see Fig. 2). Different materials, such as metals, often appear nearly identical in X-ray scans, further complicating the task of accurate object identification.
To address these challenges, we propose the X-ray Imaging-driven Detection Network (XIDNet), which leverages the unique characteristics of X-ray imaging to improve detection performance for specialized categories of prohibited items. Specifically, we introduce two key innovations: a novel X-ray-specific augmentation strategy that utilizes Poisson blending to crop instances of rare categories from their original images and seamlessly integrate them into other images, generating realistic synthetic instances and improving detection performance for categories with few samples, and an environment-aware feature enhancement algorithm that captures the spatial and semantic interactions between objects and their surroundings in X-ray images, enhancing the model’s ability to distinguish between visually similar objects by leveraging contextual relationships between target items and their environment. Together, these innovations provide a comprehensive solution to the unique challenges of X-ray security inspection.
@@
Figure 2: Preliminary experimental visualization. The left side shows the ground truth, while the right displays the predictions. We observe the difficulty in distinguishing between categories due to X-ray filtering. For example, when a hammer is present, its components are usually misclassified as a knife.
Extensive experimental evaluation demonstrates that XIDNet significantly outperforms existing SoTA methods, particularly in detecting these categories with insufficient samples. In summary, this work provides a comprehensive solution to the challenges posed by imbalanced data distributions in X-ray security screening, ultimately contributing to more reliable automated security systems. The main contributions are summarized as follows:
  • We are the first to utilize Poisson blending in X-ray-specific augmentation to generate more realistic X-ray samples, effectively addressing class imbalance and enhancing detection for tail categories with limited samples.
  • To overcome the challenge of category distinction due to color filtering in X-ray imaging, we propose a novel model-level strategy that enhances object discriminability by integrating surrounding contextual information.
@@
  • Extensive experiments show that XIDNet outperforms existing methods by a substantial margin (+17.2%), providing a robust solution for accurate prohibited item detection in real-world X-ray security inspections.
Figure 3: The framework overview of XIDNet. The proposed XIDNet consists of two key components: X-ray-specific augmentation generates synthetic data by merging X-ray objects with backgrounds, closely mimicking real-world scenarios, and contextual feature integration, which strengthens proposal features using relational context from high-confidence proposals to improve overall performance.
@@ 2.2 Long-tail Vision Tasks
Long-tail visual recognition addresses the imbalance between common (head) and rare (tail) classes in real-world datasets. Several approaches have been proposed to handle the challenge of learning effective representations for tail classes. Class rebalancing strategies like re-sampling and re-weighting aim to mitigate bias towards head classes. For example, [6] introduced an effective number of samples to create class-balanced loss weights, improving performance for underrepresented classes. Data augmentation techniques, such as [4], generate synthetic data to boost tail class instances and balance the class distribution. However, these methods may lead to overfitting or fail to capture the full diversity of tail classes.
Recent work, such as Logit Normalization (LogN) by [47], self-calibrates classified logits, similar to batch normalization, to address long-tail recognition. [45] proposed ROG, a method that reconciles object-level and global-level objectives in long-tail detection. Meta-learning and contrastive learning are also explored for improving feature learning in long-tail tasks. [15] introduced Decoupled Training, which first learns a feature extractor using cross-entropy loss and then fine-tunes it with a rebalanced classifier. [1] proposed LDAM Loss, which adjusts decision boundaries based on class frequency, increasing margins for tail classes. Additionally, contrastive learning methods like [31] align feature distributions to ensure consistency between head and tail classes, more robust to imbalance.
@@ 3 Method
In this section, we provide a detailed description of each component of the proposed XIDNet. We begin by introducing the overall framework in Sec. 3.1. Next, Sec. 3.2 covers the X-ray-specific augmentation, while Sec. 3.3 discusses the contextual feature integration. Finally, Sec. 3.4 outlines the complete network training process.
@@ 3.1 Framework Overview
Fig. 3 illustrates the proposed XIDNet framework, consisting of two main components: X-ray-specific augmentation and contextual feature integration.
@@
The X-ray-specific augmentation module addresses class imbalance, especially between head and tail categories. Instances from tail classes are cropped from their original images and treated as foreground objects. These cropped instances are then seamlessly inserted into other images using Poisson image editing, ensuring natural integration with consistent lighting and background blending. This augmentation method enhances the representation of underrepresented classes, balancing the dataset and improving the model’s ability to generalize across rare categories.
The contextual feature integration module focuses on contextual relationships between objects. By utilizing spatial and semantic correlations between the target object and surrounding items, the model can better interpret complex scenes. For example, the presence of certain objects nearby may aid in accurate classification, especially when prohibited items are often surrounded by related items. This approach reduces misclassification by leveraging spatial dependencies, improving robustness of performance in cluttered or occluded object detection under X-ray imaging.
3.2 X-ray-Specific Augmentation
Traditional data augmentation techniques like Copy-paste [10] and Mixup [42] struggle with X-ray images due to material-dependent color and texture variations. Non-metallic objects are semi-transparent or orange, while metallic ones appear darker, complicating object overlap. As shown in Fig. 1, copy-paste causes significant discrepancies, while mixup disrupts the image context. To address this, we propose an X-ray-specific augmentation strategy that maintains contextual integrity while smoothly integrating new objects, as illustrated in Fig. 3 for tail classes.
@@ 3.2 X-ray-Specific Augmentation
Consider a dataset $D=\{(I_{k},B_{k},C_{k})\}_{k=1}^{K}$, where each image $I_{k}$ contains bounding boxes $B_{k}=\{b_{1}^{k},b_{2}^{k},\dots,b_{n_{k}}^{k}\}$ and corresponding class labels $C_{k}=\{c_{1}^{k},c_{2}^{k},\dots,c_{n_{k}}^{k}\}$. In a typical long-tail distribution, the classes are divided into $C=C_{head}\cup C_{tail}$, where $C_{head}$ represents frequently occurring classes, and $C_{tail}$ represents the rarely occurring tail classes. The instances of tail classes, $N_{c}$, are much fewer than those of head classes, $N_{c^{\prime}}$, i.e., $N_{c}\ll N_{c^{\prime}}$, which limits the model’s ability to effectively generalize. For each tail class $c\in C_{tail}$, we perform the following steps:
Instance Cropping. For each image $I_{k}$ containing an instance of class $c$, the bounding box $b_{i}^{k}\in B_{k}$ corresponding to class $c$ is cropped to generate a sub-image $I_{c,i}=I_{k}[b_{i}^{k}]$. The cropped instances form the set $S_{c}=\{I_{c,i}\}$. Since the existing bounding boxes often include background elements, we manually annotate segmentation masks for precise cropping, ensuring only foreground instances are used (further details are in the supplementary materials).
Position Selection. After cropping the instances, we place them in positions corresponding to existing objects in the image, using the object labels to determine appropriate locations. This prevents placing objects in blank areas and ensures realistic overlap, as shown in Fig. 1. This method, combined with the smooth fusion process, generates varied and complex overlapping cases that improve model performance on challenging scenarios.
Poisson Fusion. We use Poisson fusion [24] for its ability to accurately model the interaction between X-ray objects and their backgrounds. Poisson fusion preserves important X-ray details, such as texture and contrast, by ensuring smooth integration of the foreground object with the background. Each cropped instance $I_{c,i}\in S_{c}$ is inserted into a new background image $I_{bg}$ using Poisson fusion to create a seamless composite image $I_{new}$, represented as:
$I_{new}=\text{PoissonFusion}(I_{bg},I_{c,i},b_{new}),$ (1)
where $b_{new}$ denotes new bounding box of inserted object.
Augmented Dataset. The newly generated images $I_{new}$, along with bounding boxes and labels, are added to the training dataset, resulting in an augmented dataset $D^{\prime}$:
$D^{\prime}=D\cup\{(I_{new},B_{new},C_{new})\},$ (2)
where $B_{new}$, $C_{new}$ are augmented bounding boxes, labels.
@@
By applying the X-ray-specific augmentation to tail class instances, we increase the effective number of training samples, reducing class imbalance and improving the model’s ability to generalize to tail classes during training.
3.3 Contextual Feature Integration
In this module, we enhance the detection process by leveraging the contextual relationships between objects and surroundings in the image. The first step is to select the top $k$ proposals based on their confidence scores. Let the set of initial proposals be $\{p_{1},p_{2},\dots,p_{N}\}$, with corresponding confidence scores $\{s_{1},s_{2},\dots,s_{N}\}$. We then select the top $k$ proposals, denoted by $P_{\text{top-k}}=\{p_{i_{1}},p_{i_{2}},\dots,p_{i_{k}}\}$, where the confidence scores satisfy $s_{i_{1}}\geq s_{i_{2}}\geq\dots\geq s_{i_{k}}$. These proposals are assumed to contain the most likely prohibited items and will serve as the basis for further enhancement.